what is the equation for; 3 times the difference of the square of a number and 1 is 72
Answer by ankor@dixie-net.com(22740)

what is the equation for; 3 times the difference of the square of a number and 1 is 72
:
Write what it says
3(x^2-1) = 72
3x^2 - 3 = 72
3x^2 = 72 + 3
3x^2 = 75
divide both sides by 3
x^2 = 25
x = 5
:
Check this
3(5^2-1) = 72